# MER 2024: Semi-Supervised Learning, Noise Robustness, and Open-Vocabulary Multimodal Emotion Recognition

### v2

@@ Abstract.
Multimodal emotion recognition is an important research topic in artificial intelligence. Over the past few decades, researchers have made remarkable progress by increasing dataset size and building more effective architectures. However, due to various reasons (such as complex environments and inaccurate labels), current systems still cannot meet the demands of practical applications. Therefore, we plan to organize a series of challenges around emotion recognition to further promote the development of this field. Last year, we launched MER2023111http://merchallenge.cn/mer2023, focusing on three topics: multi-label learning, noise robustness, and semi-supervised learning. This year, we continue to organize MER2024222https://zeroqiaoba.github.io/MER2024-website. In addition to expanding the dataset size, we introduce a new track around open-vocabulary emotion recognition. The main consideration for this track is that existing datasets often fix the label space and use majority voting to enhance annotator consistency, but this process may limit the model’s ability to describe subtle emotions. In this track, we encourage participants to generate any number of labels in any category, aiming to describe the emotional state as accurately as possible. Our baseline is based on MERTools [^3] https://github.com/zeroQiaoba/MERTools and the code is available at: https://github.com/zeroQiaoba/MERTools/tree/master/MER2024.
@@ 1. Introduction
First, it is difficult to collect large amounts of samples with emotion labels. On the one hand, the collected samples usually have no emotion (i.e. neutral) (poria2019meld, ). On the other hand, emotions have inherent subjectivity. To increase annotation consistency, researchers often hire multiple annotators and adopt majority voting (li2017reliable, ), significantly increasing labor costs. To address the problem of data sparseness, researchers revolve around unsupervised or semi-supervised learning and use unlabeled data during training. Furthermore, MERBench (lian2024merbench, ) points out the necessity of using unlabeled data from the same domain as labeled data. Therefore, we provide a large number of human-centric unlabeled videos in MER2024 and encourage participants to explore more effective unsupervised or semi-supervised learning strategies for better performance.
@@
Furthermore, emotions are subjective and ambiguous. To increase annotation consistency, existing datasets typically limit the label space to a few discrete labels. Then, they employ multiple annotators and use majority voting to select the most likely label. However, this process may cause some correct but non-candidate or non-majority labels to be ignored, resulting in inaccurate annotations. To this end, we introduce a new track around open-vocabulary emotion recognition. We encourage participants to generate any number of labels in any category, trying to describe emotions accurately.
@@ 2. Challenge Dataset
MER2024 is an extended version of MER2023 (lian2023mer, ). Figure 1 shows the dataset construction pipeline. Specifically, MER2023 contains four subsets: Train&\&&Val, MER-MULTI, MER-NOISE, and MER-SEMI. In the last subset, besides labeled samples, it also contains a large amount of unlabeled data. In MER2024, we merge all labeled samples and get the updated Train&\&&Val. Meanwhile, we collect more unlabeled data and select some unlabeled samples for annotation, getting MER-SEMI and MER-NOISE. For MER-OV, we select 332 samples from Train&\&&Val and rely on LLMs to extract open-vocabulary labels from emotion-related descriptions (see Section 2.3 for more details). Table 1 provides statistics for MER2023 and MER2024.
@@ 2.3. MER-OV
MER-OV focuses on open-vocabulary emotion recognition. Unlike MER-SEMI and MER-NOISE which only predict the most likely emotion within a fixed label space, MER-OV needs to predict any number of emotions in any category, trying to describe the emotional state as accurately as possible. Previously, we made an initial attempt on this task (lian2023explainable, ). Specifically, we first annotate emotion-related acoustic and visual clues. Then, we rely on the LLMs’ reasoning ability and use these clues to disambiguate video subtitles. This process can generate descriptions containing more subtle emotional labels. Finally, we use ChatGPT to extract all labels using the following prompt:
Please assume the role of an expert in the field of emotions. We provide clues that may be related to the emotions of the characters. Based on the provided clues, please identify the emotional states of the main characters. Please separate different emotional categories with commas and output only the clearly identifiable emotional categories in list format. If none are identified, please output an empty list.
Due to the high annotation cost, we only select 332 samples for annotation (lian2023explainable, ). Through this process, we can generate an average of 2.92 labels per sample. We encourage participants to use MLLMs directly or to further conduct instruction fine-tuning with our annotated samples to solve this task.
@@
Audio LLM Subtitle content of the audio: {subtitle}; As an expert in the field of emotions, please focus on the acoustic information and subtitle content in the audio to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual in the audio.
Video LLM Subtitle content of the video: {subtitle}; As an expert in the field of emotions, please focus on the facial expressions, body movements, environment, subtitle content, etc., in the video to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual in the video.
Audio-Video LLM Subtitle content of the video: {subtitle} As an expert in the field of emotions, please focus on the facial expressions, body movements, environment, acoustic information, subtitle content, etc., in the video to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual in the video.
@@ 2.4. Challenge Protocol
To download the dataset, participants should fill out an End User License Agreement (EULA) [^8] https://drive.google.com/file/d/1cXNfKHyJzVXg_7kWSf_nVKtsxIZVa517/view?usp=sharing. It requires participants to use this dataset only for academic research and not to edit or upload it to the Internet. For MER-SEMI and MER-NOISE, each team should submit the most likely discrete label among the 6 candidate labels (i.e., worried, happy, neutral, angry, surprised, and sad). For MER-OV, each team can submit any number of labels in any category. For all tracks, participants should predict results for all 115,595 unlabeled data, although we only evaluate a small subset. This requires participants to focus on the generalization ability and develop systems that do not require a large amount of inference time. Additionally, it is worth noting that participants cannot manually annotate the samples in MER2024. We will ask them to submit their code for further checking. Finally, each team should submit a paper describing their method, otherwise, the submitted results will not be included in the final leaderboard.
@@ 3.2. MER-OV
Open-vocabulary emotion recognition is a new task that lacks large-scale training data. Considering that MLLMs can handle various multimodal tasks, we choose MLLMs as the baselines. Meanwhile, temporal information is important for understanding the emotions. Therefore, we select MLLMs that support at least audio or video input, including video LLMs (such as VideoChat (li2023videochat, )), audio LLMs (such as SALMONN (tang2023salmonn, )), and audio-video LLMs (such as PandaGPT (su2023pandagpt, )). In practice, we first use the prompt in Table 2 to generate emotion-relevant descriptions that contain multi-faceted and multi-modal clues. During inference, we add subtitles as additional input. Then, we use the prompt in Section 2.3 to extract all emotion labels in an open-vocabulary manner.
@@ 3.4. Evaluation Metrics
For MER-SEMI and MER-NOISE, we choose two widely used metrics in emotion recognition: accuracy and weighted average F-score (WAF) (lian2021ctnet, ). Considering the inherent class imbalance, we choose WAF as the final ranking. For MER-OV, we draw on our previous work (lian2023explainable, ), in which we extend traditional classification metrics (i.e., accuracy and recall) and define set-level metrics. Specifically, assume that the true label set is $\{y_{i}\}_{i=1}^{M}$ and the predicted label set is $\{\hat{y}_{i}\}_{i=1}^{N}$, where $M$ and $N$ are the number of labels. Considering that there may be labels with different expressions but the same meaning, we first use ChatGPT to group all labels:
Please assume the role of an expert in the field of emotions. We provide a set of emotions. Please group the emotions, with each group containing synonyms or consistent emotional terms. Directly output the results, with each group in list format.
@@
To obtain more accurate open-vocabulary labels, we first extract labels using the approach in Section 2.3. Then, we merge these labels with the predictions submitted by participants. Finally, we manually check all labels to obtain the final ground truth.
@@ 4.1. MER-SEMI and MER-NOISE
Table 5. Performance of different methods on MER-OV (%). These results are drawn from our previous work (lian2023explainable, ). The values in the gray column are used for the final ranking. Here, “L”, “V”, and “A” indicate whether lexical, visual, and acoustic information is used in the emotion recognition process.
@@
Empty $\times$ $\times$ $\times$ 0.0 0.0 0.0
Random $\times$ $\times$ $\times$ 19.0 24.1 13.8
Valley (luo2023valley, ) $\surd$ $\surd$ $\times$ 11.1 12.7 9.6
Otter (li2023otter, ) $\surd$ $\surd$ $\times$ 24.5 27.6 21.4
PandaGPT (su2023pandagpt, ) $\surd$ $\surd$ $\surd$ 25.3 26.5 24.2
Video-LLaMA (zhang2023video, ) $\surd$ $\surd$ $\times$ 28.4 27.3 29.4
VideoChat2 (li2024mvbench, ) $\surd$ $\surd$ $\times$ 32.9 36.8 28.9
Video-ChatGPT (maaz2023video, ) $\surd$ $\surd$ $\times$ 35.3 39.0 31.7
VideoChat (li2023videochat, ) $\surd$ $\surd$ $\times$ 35.9 34.3 37.5
SALMONN (tang2023salmonn, ) $\surd$ $\times$ $\surd$ 36.0 34.3 37.8
mPLUG-Owl (ye2023mplugowl, ) $\surd$ $\surd$ $\times$ 39.8 38.8 40.7
Qwen-Audio (chu2023qwen, ) $\surd$ $\times$ $\surd$ 46.0 54.0 38.0
GPT-4V (openai2023gpt4v, ) $\surd$ $\surd$ $\times$ 56.0 53.2 58.9
@@ 4.2. MER-OV
In this section, we report the performance of different methods on this track. Besides MLLMs, we introduce two heuristic baselines: Empty and Random. For the former, we do not assign any label. For the latter, we randomly select a label from six candidate categories (i.e., worried, happy, neutral, angry, surprised, and sad). Considering that the emotion labels extracted in Section 2.3 are relatively accurate, we use them as the ground truth. Table 5 shows the performance of different methods. Experimental results demonstrate that MLLMs generally outperform heuristic baselines, showcasing that MLLMs can handle this task to some extent. However, these models still cannot address this task perfectly. Meanwhile, audio LLMs generally perform better than video LLMs. The reason is that the samples in our dataset mainly convey emotions through the acoustic modality, which is consistent with the findings in Table 3.

### v3

@@ Abstract.
Multimodal emotion recognition is an important research topic in artificial intelligence. Over the past few decades, researchers have made remarkable progress by increasing dataset size and building more effective architectures. However, due to various reasons (such as complex environments and inaccurate annotations), current systems are hard to meet the demands of practical applications. Therefore, we organize a series of challenges around emotion recognition to further promote the development of this area. Last year, we launched MER2023111http://merchallenge.cn/mer2023, focusing on three topics: multi-label learning, noise robustness, and semi-supervised learning. This year, we continue to organize MER2024222https://zeroqiaoba.github.io/MER2024-website. In addition to expanding the dataset size, we introduce a new track around open-vocabulary emotion recognition. The main consideration for this track is that existing datasets often fix the label space and use majority voting to enhance annotator consistency, but this process may limit the model’s ability to describe subtle emotions. In this track, we encourage participants to generate any number of labels in any category, aiming to describe the emotional state as accurately as possible. Our baseline is based on MERTools [^3] https://github.com/zeroQiaoba/MERTools and code is available at: https://github.com/zeroQiaoba/MERTools/tree/master/MER2024.
@@ 1. Introduction
First, it is difficult to collect large amounts of samples with emotion labels. On the one hand, the collected samples usually have no emotion (i.e. neutral) (poria2019meld, ). On the other hand, emotion annotation relies on subjective judgment. To increase annotation consistency, researchers often hire multiple annotators and adopt majority voting (li2017reliable, ), significantly increasing labor costs. To address the problem of data sparseness, researchers revolve around unsupervised or semi-supervised learning and use unlabeled data during training. During unsupervised or semi-supervised training, MERBench (lian2024merbench, ) points out the necessity of using unlabeled data from the same domain as labeled data. Therefore, we provide a large number of human-centric unlabeled videos in MER2024 and encourage participants to explore more effective unsupervised or semi-supervised learning strategies for better performance.
@@
Furthermore, emotion annotation relies on subjective judgment. To increase annotation consistency, existing datasets typically limit the label space to a few discrete labels. Then, they employ multiple annotators and use majority voting to select the most likely label. However, this process may cause some correct but non-candidate or non-majority labels to be ignored, resulting in inaccurate annotations. To this end, we introduce a new track around open-vocabulary emotion recognition. We encourage participants to generate any number of labels in any category, trying to describe emotions accurately.
@@ 2. Challenge Dataset
MER2024 is an extended version of MER2023 (lian2023mer, ). Figure 1 shows the dataset construction pipeline. Specifically, MER2023 contains four subsets: Train&\&&Val, MER-MULTI, MER-NOISE, and MER-SEMI. In the last subset, besides labeled samples, it also contains a large amount of unlabeled data. In MER2024, we merge all labeled samples and get the updated Train&\&&Val. Meanwhile, we collect more unlabeled data and select some unlabeled samples for annotation, getting MER-SEMI and MER-NOISE. For MER-OV, we select 332 samples from Train&\&&Val and provide open-vocabulary labels. Table 1 provides statistics for MER2023 and MER2024.
@@ 2.3. MER-OV
MER-OV focuses on open-vocabulary emotion recognition. Unlike MER-SEMI and MER-NOISE which only predict the most likely emotion within a fixed label space, MER-OV needs to predict any number of emotions in any category, trying to describe the emotional state as accurately as possible. Previously, we made an initial attempt on this task (lian2023explainable, ). Specifically, we first annotate emotion-related acoustic and visual clues. Then, we rely on the LLMs’ reasoning ability and use these clues to disambiguate video subtitles. This process can generate descriptions containing more subtle emotional labels. Finally, we use GPT-3.5 (“gpt-3.5-turbo-16k-0613”) to extract all labels using the following prompt:
Please assume the role of an expert in the field of emotions. We provide clues that may be related to the emotions of the characters. Based on the provided clues, please identify the emotional states of the main characters. Please separate different emotional categories with commas and output only the clearly identifiable emotional categories in a list format. If none are identified, please output an empty list.
Due to the high annotation cost, we only select 332 samples for annotation (lian2023explainable, ). Through this process, we can generate an average of 3 labels per sample. Then, we further conduct manual checks and get the final open-vocabulary labels $Y_{gt}$. We encourage participants to use MLLMs directly or to further conduct instruction fine-tuning with our annotated samples to solve this task.
@@
Audio LLM As an expert in the field of emotions, please focus on the acoustic information in the audio to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual.
Video LLM As an expert in the field of emotions, please focus on the facial expressions, body movements, environment, etc., in the video to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual in the video.
Audio-Video LLM As an expert in the field of emotions, please focus on the facial expressions, body movements, environment, acoustic information, etc., in the video to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual in the video.
@@ 2.4. Challenge Protocol
To download the dataset, participants should fill out an End User License Agreement (EULA) [^8] https://drive.google.com/file/d/1cXNfKHyJzVXg_7kWSf_nVKtsxIZVa517/view?usp=sharing. It requires participants to use this dataset only for academic research and not to edit or upload it to the Internet. For MER-SEMI and MER-NOISE, each team should submit the most likely discrete label among the 6 candidate labels (i.e., worried, happy, neutral, angry, surprised, and sad). For MER-OV, each team can submit any number of labels in any category. Meanwhile, we require that participants cannot use closed-source models (such as GPT or Claude) for MER-OV. For all tracks, participants should predict results for all 115,595 unlabeled data, although we only evaluate a small subset. This requires participants to focus on the generalization ability and develop systems that do not require a large amount of inference time. Additionally, it is worth noting that participants cannot manually annotate the samples in MER2024. We will ask them to submit their code for further checking. Finally, each team should submit a paper describing their method.
@@ 3.2. MER-OV
Open-vocabulary emotion recognition is a new task that lacks large-scale data. Considering that MLLMs can handle various multimodal tasks, we choose MLLMs as the baselines. Meanwhile, temporal information is important for understanding the emotions. Therefore, we select MLLMs that support at least audio or video input, including video LLMs (such as VideoChat (li2023videochat, )), audio LLMs (such as SALMONN (tang2023salmonn, )), and audio-video LLMs (such as PandaGPT (su2023pandagpt, )). In practice, we first use the prompt in Table 2 to generate emotion-relevant descriptions that contain multi-faceted and multi-modal clues. Then, considering that the subtitle contains important clues for emotion recognition, we use the extracted clues to disambiguate the subtitle. Finally, we use the prompt in Section 2.3 to extract all emotion labels in an open-vocabulary manner.
@@ 3.4. Evaluation Metrics
For MER-SEMI and MER-NOISE, we choose two widely used metrics in emotion recognition: accuracy and weighted average F-score (WAF) (lian2021ctnet, ). Considering the inherent class imbalance, we choose WAF as the final ranking. For MER-OV, we draw on our previous work (lian2023explainable, ), in which we extend traditional classification metrics (i.e., accuracy and recall) and define set-level metrics. Specifically, assume that the true label set is $\{y_{i}\}_{i=1}^{M}$ and the predicted label set is $\{\hat{y}_{i}\}_{i=1}^{N}$, where $M$ and $N$ are the number of labels. Considering that there may be labels with different expressions but the same meaning, we first use GPT-3.5 to group all labels:
Please assume the role of an expert in the field of emotions. We provide a set of emotions. Please group the emotions, with each group containing emotions with the same meaning. Directly output the results. The output format should be a list containing multiple lists.
@@ 4.1. MER-SEMI and MER-NOISE
Table 5. Performance of different methods on MER-OV (%). These results are drawn from our previous work (lian2023explainable, ). The values in the gray column are used for the final ranking. Here, “L”, “V”, and “A” indicate whether lexical, visual, and acoustic information is used in emotion recognition.
@@
Empty $\times$ $\times$ $\times$ 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00
Random $\times$ $\times$ $\times$ 19.13$\pm$0.06 24.85$\pm$0.15 13.42$\pm$0.04
Qwen-Audio (chu2023qwen, ) $\surd$ $\times$ $\surd$ 40.23$\pm$0.09 49.42$\pm$0.18 31.04$\pm$0.00
OneLLM (han2023onellm, ) $\surd$ $\times$ $\surd$ 43.04$\pm$0.06 45.92$\pm$0.05 40.15$\pm$0.06
Otter (li2023otter, ) $\surd$ $\surd$ $\times$ 44.40$\pm$0.09 50.71$\pm$0.10 38.09$\pm$0.09
VideoChat (li2023videochat, ) $\surd$ $\surd$ $\times$ 45.70$\pm$0.09 42.90$\pm$0.27 48.49$\pm$0.10
Video-LLaMA (zhang2023video, ) $\surd$ $\surd$ $\times$ 44.74$\pm$0.14 44.14$\pm$0.13 45.34$\pm$0.15
PandaGPT (su2023pandagpt, ) $\surd$ $\surd$ $\surd$ 46.21$\pm$0.17 50.03$\pm$0.01 42.38$\pm$0.33
SALMONN (tang2023salmonn, ) $\surd$ $\times$ $\surd$ 48.06$\pm$0.04 50.20$\pm$0.04 45.92$\pm$0.04
Video-LLaVA (lin2023video, ) $\surd$ $\surd$ $\times$ 47.12$\pm$0.15 48.58$\pm$0.02 45.66$\pm$0.29
VideoChat2 (li2024mvbench, ) $\surd$ $\surd$ $\times$ 49.60$\pm$0.28 54.72$\pm$0.41 44.47$\pm$0.15
OneLLM (han2023onellm, ) $\surd$ $\surd$ $\times$ 50.99$\pm$0.08 55.93$\pm$0.09 46.06$\pm$0.06
LLaMA-VID (li2023llama, ) $\surd$ $\surd$ $\times$ 51.29$\pm$0.09 52.71$\pm$0.18 49.87$\pm$0.00
mPLUG-Owl (ye2023mplug, ) $\surd$ $\surd$ $\times$ 52.79$\pm$0.13 54.54$\pm$0.13 51.04$\pm$0.13
Video-ChatGPT (maaz2023video, ) $\surd$ $\surd$ $\times$ 50.73$\pm$0.06 54.03$\pm$0.04 47.44$\pm$0.07
Chat-UniVi (jin2023chat, ) $\surd$ $\surd$ $\times$ 53.09$\pm$0.01 53.68$\pm$0.00 52.50$\pm$0.02
GPT-4V (openai2023gpt4v, ) $\surd$ $\surd$ $\times$ 56.69$\pm$0.04 48.52$\pm$0.07 64.86$\pm$0.00
@@ 4.2. MER-OV
In this section, we report the performance of different methods on this track. Besides MLLMs, we introduce two heuristic baselines: Empty and Random. For the former, we do not assign any label. For the latter, we randomly select a label from six candidate categories (i.e., worried, happy, neutral, angry, surprised, and sad). Table 5 shows the performance of different methods. Experimental results demonstrate that MLLMs generally outperform heuristic baselines, showcasing that MLLMs can handle this task to some extent. However, these models still cannot address this task perfectly, indicating the difficulty of this task.